MARS DESERT RESEARCH STATION

Daily Summary Report – December 25th

Daily Summary Report

SOL: 07

Person filling out Report: Alison Gibson, Commander

Summary Title: A Windy Christmas

Mission Status: Ready for a productive week

Sol Activity Summary: We started the day with a Christmas breakfast feast, did our Yankee Swap, played games, and ended the afternoon with a movie.

Look Ahead Plan: EVA to Candor Chasma

Anomalies in work:  We had an intense wind storm last night and the cover blew off the hole in the top of the Hab. It was reinstalled as best as possible, but may need future improvements.

Weather: Partly cloudy, High 45F, Low 33F, wind gusts up to 33 mph

Crew Physical Status: Fully functional

EVA: Our EVA was postponed until tomorrow to avoid strong winds and celebrate the holiday

Reports to be filed: Sol Summary, Operations Report, Science Report

Support Requested: None

Daily Summary Report – December 24th

Daily Summary Report
SOL: 06
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Christmas Eve
Mission Status: Research moving along, but slowly.
Sol Activity Summary: Slow day. Brought greenhab plants inside.
Look Ahead Plan: EVA to Candor Chasma tomorrow
Anomalies in work: Nothing
Weather: High 33F, Low 48F, wind avg 6.5mph, gust 7.4mph, precip
0.02″, grey cloudy skies
Crew Physical Status: Active. Full crew functional.
EVA: Postponed due to prior precipitation. Too muddy.
Reports to be filed:
– Sol Summary
– Science Reports
– 6-8 Photos
– EVA Plan

Sol Summary – December 23rd

Daily Summary Report
SOL: 05
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Week One Nearly Complete
Mission Status: Research moving along, but slowly.
Sol Activity Summary: Second EVA to dinosaur quarry, minor greenhab concerns
Look Ahead Plan: Christmas is coming up!
Anomalies in work: None
Weather: High 37F, Low 31F, wind avg 3mph, gust 4.5mph, precip 0.04″,
grey cloudy skies
Crew Physical Status: Active. Full crew functional.
EVA: Crew B to dinosaur quarry
Reports to be filed:
– Sol Summary
– Journalist’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested
– None

Sol Summary – December 22nd

Daily Summary Report
SOL: 04
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Snowday
Mission Status: Active. Full crew functional.
Sol Activity Summary: Snow on ground, greenhouse work continues, EVA postponed
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: Nothing
Weather: High 35F, Low 29F, wind avg 1.5mpg, gust N/A, humidity
35-98%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Postponed until tomorrow
Reports to be filed:
– Sol Summary
– Journalist’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested
– None

Sol Summary – December 21st

SOL: 03
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Longer EVA, beginning greenhab setup
Mission Status: Active. Full crew functional.
Sol Activity Summary: EVA’d to Dinosaur Quarry, set up first conveyor
stage in greenhab, planted first 12 plants.
Look Ahead Plan: EVA tomorrow to follow up on initial Quarry recon
Anomalies in work: back side of helmet cracked slightly on rock.
Repaired with superglue and zip ties. Not majorly concerning, but
helmet henceforth moved to reserve use only.
Weather: High 42F, Low 18F,  wind avg 3.5mph, gust N/A, humidity
17-41%, grey cloudy skies
Crew Physical Status: Slightly stir crazy, but ok.
EVA: Anselm, Geoff, and Alison to Dinosaur Quarry
Reports to be filed:
– Sol Summary
– Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested:
– Wifi still intermittent.

Sol Summary – December 20th

Daily Summary Report
SOL: 02
Person filling out Report: Anselm Wiercioch, XO
Summary Title: First Recon
Mission Status: Crew is alive and well
Sol Activity Summary: Received water refil, went on first EVA
Look Ahead Plan: Planning longer recon EVA tomorrow, considering
sealing water connection further.
Anomalies in work: None significant.
Weather: High 51F, Low 10F,  wind avg 11.9mph, gust 12.3mph, humidity
12-45%, clear and sunny skies.
Crew Physical Status: Less nervous, less thirsty. Generally OK.
EVA: Scouted local area around hab. Explored northwestern ridge and
stream to the north of hab.
Reports to be filed:
– Sol Summary
– Journalist/Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested:
– Will be keeping an eye on internet connectivity, but generally OK
for the time being.

Sol Summary – December 19th

Daily Report:

SOL: 01
Person filling out Report: Anselm Wiercioch, XO
Summary Title: Waking up on Mars.
Mission Status: Active. Full crew functional.
Sol Activity Summary: First day on Mars. Set up new water tank, hot water heater, greenhab thermometer.
Look Ahead Plan: Prepping for first exploratory EVA tomorrow, testing new water connection, beginning greenhab setup.
Anomalies in work: Hole in water line, propane smell around furnace.
Weather: sunrise 7:36, sunset 17:05, high 71F, low 28F, generally sunny and clear skies.
Crew Physical Status: Nervous, thirsty.
EVA: None.
Reports to be filed:
– Sol Summary (herein)
– Journalist/Commander’s Report
– Science Reports
– 6-8 Photos
– EVA Plan
– Operations Report
Support Requested:
– Should we be expecting a water refill soon? The new tank was moved into position, but we’re still running on the ~200 gallons or so that were already on the trailer. Not sure how long that will last.
– We can smell propane near the furnace. Very faint and not immediately concerning, but will be keeping an eye on it. Maybe have someone on call in case it becomes more noticeable.